OpenStudy (anonymous):

(cot(x)-tan(x))/(sin(x)cos(x)) = csc^2(x)-sec^(2)x

OpenStudy (anonymous):

i need to prove it

OpenStudy (campbell_st):

rewrite cot ans tan in terms of sin and cos \[\frac{\frac{\cos(x)}{\sin(x)} - \frac{\sin(x)}{\cos(x)}}{\sin(x)\cos(x)} = \frac{\frac{\cos^2(x) - \sin^2(x)}{\sin(x)\cos(x)}}{\sin(x)\cos(x)} \] this can be simplified to \[\frac{\cos^2(x) - \sin^2(x)}{\sin^2(x)\cos^2(x)} \] now just split the fraction \[\frac{\cos^2(x)}{\sin^2(x)\cos^2(x)} - \frac{\sin^2(x)}{\sin^2(x)\cos^2(x)}\] just remove the common factors for the answer.

OpenStudy (anonymous):

cot x = cosx/sinx tan x = sinx/cosx Take Left side first (cosx/sinx - sinx/cosx) (1/sinx cosx) Now take common denominator (cos^2 x - sin^2 x)/sinx cosx (1/sinx cosx) Multiply the common denominator with the denominator give in the equation (cos^ 2 x - sin^2 x) / sin^2 x cos^ 2x Separate both terms (cos^2 x/cos^2 x sin^2 x) - (sin^2 x/ sin^2 x cos^2 x) Cancel out cosine and sines 1/sin^2 x - 1/cos^2 x csc^2 x - sec^2 x which is equal to the right side